### v2

# Beyond Early-Token Bias: Model-Specific and Language-Specific Position Effects in Multilingual LLMs

Abstract
Large Language Models (LLMs) exhibit position bias—a systematic tendency to neglect information at specific context positions. However, the patterns of position bias behavior, depending on the language or model, remain unexplored. We present a multilingual study across five typologically distinct languages (English, Russian, German, Hindi, and Vietnamese) and five model architectures, examining how position bias interacts with prompt strategies and affects output entropy. Our key findings are: (1) Position bias is primarily model-driven, yet exhibits language-specific variations. For instance, Qwen2.5-7B-Instruct and DeepSeek 7B Chat consistently favors late positions, challenging established assumptions of a universal early-token bias in LLMs. (2) Explicitly instructing the model that "the context is relevant to the query" unexpectedly reduces accuracy across languages, undermining common prompt-engineering practices. (3) While the largest accuracy drop occurs when relevant information is placed in the middle of the context, this is not explicitly reflected by a corresponding peak in output entropy.
1 Introduction
Many recent applications based on large language models (LLMs) require support for long-context processing [^1] Such applications may include Retrieval-Augmented Generation, autonomous agents, scientific research, customer support, among others.. However, developing new training strategies to accommodate longer contexts is insufficient, as new challenges continue to arise. One notable issue is position bias – the systematic neglect of information located at specific positions, typically in the middle of the context (Baker et al., 2024)
While position bias has been well-documented in English-centric studies (Zhang et al., 2024b; Baker et al., 2024), its manifestation in multilingual contexts and across-various architectures remains underexplored. Furthermore, existing bias mitigation strategies (Peysakhovich & Lerer, 2023; Zhao et al., 2021; Zhang et al., 2024a; Guo & Vosoughi, 2024) have predominantly been evaluated on English-language datasets. As discussed in Section 2, lexical and morphological variations across languages necessitate careful consideration in multilingual modeling. As a step toward adapting multilingual large language models for real-world long-context applications, this work seeks to address the following research questions:
Q1: Is position bias primarily a model-driven phenomenon, or do language-specific patterns emerge due to lexical, morphological, and syntactic differences (Ghosh et al., 2024b)?
Q2: Do prompt-based strategies (e.g., explicit position guidance (Zhang et al., 2024a)) effectively mitigate bias across languages? Can we focus model attention via prompting, and would it improve model performance?
Q3: Finally, we are interested in a deeper understanding of how position bias affects model generation, and thus we perform formal and empirical analysis of how position bias affects entropy of the output distribution.
Contributions. We present a multilingual analysis of position bias across five typologically diverse languages (English, German, Russian, Hindi, and Vietnamese) and five model architectures. Our key contributions are:
  • We show that position bias is predominantly model-driven, yet exhibits significant language-specific variations. For instance, models like Qwen2.5-7B-Instruct (Qwen et al., 2025) and DeepSeek-7B-Chat (DeepSeek-AI, 2024) display a strong late-position bias, contradicting prior claims of an inherent early-token preference in LLMs (Wu et al., 2025; Barbero et al., 2025). In contrast, Llama3.1-8B-Instruct (Touvron et al., 2023) prioritizes early positions. We speculate these differences stem from variations in training data and model architecture.
  • We demonstrate that explicitly instructing models about correct context placement (e.g., “the correct context has label 1”) consistently degrades accuracy across all languages. This finding challenges established prompt-engineering practices (Zhang et al., 2024a).
  • Through an empirical analysis of output entropy, we reveal a counterintuitive dynamic: while model accuracy is lowest when relevant context is in the middle, this performance drop is not accompanied by a corresponding peak in entropy. This indicates a disconnect between positional disadvantage and the model’s expressed uncertainty.
These findings yield the following practical implications:
  1. 1.
    Chain-of-Thought and other reasoning strategies that rely on explicit positional guidance demand careful reconsideration, as our results show they can cause consistent performance degradation.
  2. 2.
    Retrieval-augmented generation (RAG) systems must account for specific model and language characteristics. Strategies that reorder context based on the assumption that models prioritize recent or initial tokens (see Section 2) may be ineffective or detrimental for models with a late-position bias.
  3. 3.
    The observed entropy dynamics complicate uncertainty-based bias mitigation techniques (Duan et al., 2024) and are crucial for developing effective uncertainty quantification strategies in RAG pipelines.
The code for our experiments is available in the GitHub repository.
2 Related Work
What causes position bias? Prior work identifies multiple contributing factors. (Zhang et al., 2024b) attribute position bias to U-shaped attention patterns in transformers, which prioritize extremal positions and degrade performance for mid-context evidence. Theoretical and empirical studies further demonstrate that transformers attention is inherently biased toward earlier tokens (Wu et al., 2025; Barbero et al., 2025). (Wu et al., 2025) explains position bias due to "causal masking inherently biases attention toward earlier positions, as tokens in deeper layers attend to increasingly contextualized representations of earlier tokens". Our results reveal exceptions where later positions are desired, highlighting the complexity of the problem. Training data biases, such as serial-position effects in corpora, shape how models prioritize sequence positions (Wu et al., 2025; Guo & Vosoughi, 2024).
Interplay of Culture, Language, and Model Design. The way we perceive the world is influenced not only by our culture (Oyserman & Lee, 2008) but also by the language we speak (Boroditsky et al., 2003). The latter point is particularly relevant for LLMs, since they are trained on specific languages. Recent studies have shown that multilingual LLMs often initiate their "thinking process" in English, pivoting to the prompt’s original language in the middle layers (Zhang et al., 2024b; Peysakhovich & Lerer, 2023). These models exhibit lower lexical naturalness in non-English languages, with the naturalness gap being more pronounced for languages structurally distant from English (Guo et al., 2024). While the volume of training data plays a crucial role (Arnett & Bergen, 2025), linguistic complexity — including lexical and morphological variations across languages — must also be considered (Ghosh et al., 2024b; Dang et al., 2024; Ismayilzada et al., 2025). Additionally, architectural design choices affect languages in different ways; for instance, removing positional encoding from language models would most degrade performance in languages with limited morphological systems (Ghosh et al., 2024b).
At the same time, most of bias mitigation approaches evaluate their performance in English (Zhang et al., 2024b; Peysakhovich & Lerer, 2023; Zhang et al., 2024a; Yu et al., 2024b; Wang et al., 2024b). These approaches fall into two categories: prompt-based techniques and architectural interventions. Architectural methods—such as positional encodings, alternative masking schemes, and calibration mechanisms—address root causes but often require retraining and introduce computational overhead (Zhang et al., 2024b; Wu et al., 2025; Zhao et al., 2021). Prompt-based strategies, including query-aware contextualization and recency prompting, aim to redirect attention dynamically (Peysakhovich & Lerer, 2023; Wang et al., 2024b; Yu et al., 2024b). We focus on prompt-based strategies, using as a starting point the work done by (Zhang et al., 2024a). The authors studied whether a model can improve its performance when given explicit placement of the correct answer. They used two types of instructions — relative and absolute — and found that models lack relative awareness and that implicit information about absolute placement of the correct prompt marginally improves model performance. However, the authors did not study a scenario when the placement prompt is absent at all.
Practical considerations. Position bias affects Chain of Thought Strategies (CoT). CoT struggles with position bias even when reasoning steps are correct, models often fail to retrieve evidence from middle positions (Zhang et al., 2024b). In (Zhang et al., 2024b; Yu et al., 2024b), authors analyzed error propagation in multi-hop reasoning. In RAG systems, one of the mitigating strategies is context ordering  (Wang et al., 2024a; Alessio et al., 2024; Jin et al., 2024) (we discuss these approaches in Appendix A). While conventional approaches often assume a monotonic relationship between document position and attention (e.g., prioritizing the first/last positions), our analysis reveals that position bias may exhibit language-specific patterns and is not always maximized at early tokens. This observation challenges assumptions in methods like Long-Context RAG, which rely on fixed position prioritization, and highlights the need for language-adaptive reordering.
While predictive entropy is widely used to quantify model uncertainty (Huang et al., 2024; Sychev et al., 2025), its relationship with position bias remains unexplored. (Duan et al., 2024) note that uncertainty estimates can be token-biased, but how position bias interacts with uncertainty dynamics is unclear.
3 Methods
3.1 Position bias formalization
To evaluate the effect of position bias, we consider a question-answering task. For each question $Q$, we assume the existence of a ground truth answer $A$ and a set of $N$ contexts $\text{ctx}_{1},\dots,\text{ctx}_{N}$. In our experimental setup, exactly one context is relevant to each question, while the remaining N1N-1 contexts are randomly sampled from the dataset. The response of the model, denoted as Model, using a prompt function Prompt, to the question $Q$ when the relevant context is placed at position $i$ is given by: $R_{i}=\text{Model}(\text{Prompt}(Q,\text{ctx}_{1},\dots,\text{ctx}_{N},i))$. We define that the model exhibits position bias on dataset $D$ toward position $i$ over position $j$ if the expected Accuracy when the relevant context is at position $i$ exceeds that at position $j$:
$\mathop{\mathbb{E}}_{(Q,A)\sim D}\left[\text{Acc}(A,R_{i})\right]>\mathop{\mathbb{E}}_{(Q,A)\sim D}\left[\text{Acc}(A,R_{j})\right]$
3.2 Context Placement Strategies
Figure 1: Experiment Structure. For each question, the relevant context is placed in one of three positions: top, middle, or bottom of the context list. Each context is assigned a binary score (0 or 1), indicating its relevance to the question. Three scoring strategies are evaluated: Aligned, where the relevant context receives a score of 1 and all others are assigned 0; All Zero, in which every context, including the relevant one, is scored 0; No Scores, where relevance scores are entirely omitted. This experimental design assesses the influence of scoring mechanisms and answer positioning on model performance under varying degrees of contextual guidance.
This experimental series examines whether explicit information regarding the relevance of contexts to a given query can improve performance. We investigate a practical scenario in which context relevance can be quantified (e.g., through cosine similarity between context and question embeddings) and assess the impact of integrating relevance scores into prompts on context selection. The experimental framework is illustrated in Figure 1.
Scoring Strategies. Three distinct scoring strategies are evaluated. The Aligned strategy assigns a relevance score of $rs=1$ to contextually relevant information and $rs=0$ to all other contexts. The All Zero strategy assigns $rs=0$ to all contexts, including those that are relevant, to test the hypothesis that intentional mislabeling will degrade model performance if relevance scores influence context selection. The No Scores strategy omits relevance scores entirely from the prompts to assess whether explicit scoring contributes to model efficacy. In the experimental setup, relevance scores $rs_{1},\ldots,rs_{N}$ were incorporated into the prompt where applicable. As illustrated in Figure 1, three positions for relevant contexts are considered: TOP (first), MIDDLE ($N/2$), and BOTTOM (last).
Context Volume. Prior research has shown that longer contexts can exacerbate position bias (Baker et al., 2024; Peysakhovich & Lerer, 2023). In our work, we default to five contexts based on preliminary experiments. We vary the number of contexts, setting $N\in\{5,10,15\}$ , and find that while increasing N does degrade performance for some models, the effect is relatively modest. Given limited computational resources, we therefore opted to use five contexts in our main experiments. Preliminary results are provided in Appendix G.
3.3 Average Predictive Entropy
We adopt the token-wise entropy framework introduced by (Lu et al., 2022), normalized by the total number of tokens, to quantify uncertainty in model responses (Lu et al., 2022; Wang et al., 2025). Let $x$ represent the input prompt and $s=\{z_{1},z_{2},\dots,z_{n}\}$ denote a generated completion sequence of $n$ tokens. For a given large language model (LLM), the conditional probability of generating the $i$-th token $z_{i}$, given the preceding tokens $s_{<i}=\{z_{1},\dots,z_{i-1}\}$ and the prompt $x$, is denoted as $p(z_{i}\mid s_{<i},x)$ for $1\leq i\leq n$. The average predictive entropy1 (denoted as $\textit{PE}_{\text{avg}}$) is defined as:
$\textit{PE}_{\text{avg}}(s,x)=-\frac{1}{n}\log p(s\mid x)=\frac{1}{n}\sum_{i=1}^{n}-\log p(z_{i}\mid s_{<i},x).$ (1)
This formulation computes the average uncertainty per token by decomposing the joint probability $p(s\mid x)$ into a product of conditional probabilities using the chain rule. The normalization by $n$ ensures comparability across sequences of varying lengths, consistent with the interpretation of entropy as a measure of "average uncertainty".
4 Experiment Set-Up
4.1 Datasets
In this study, we employed three open-ended question-answering datasets encompassing five languages characterized by divergent syntactic structures and semantic distributions. The statistics for these datasets are summarized in Table 1, with a comprehensive description provided in Appendix B. These datasets were selected for two principal reasons: (1) they are well-established within the research community, ensuring familiarity and reproducibility; (2) the context lengths for question-answer pairs are sufficiently concise (less than 4096 characters), allowing multiple instances to be included in a single prompt without exceeding the maximum sequence length constraints of the language model.
Table 1: Summary of datasets utilized in this study, categorized by language, with corresponding sources and associated question-answer pairs sizes
Language Source Size
English SQuAD2.0 (Rajpurkar et al., 2018) 150k
Russian MTS-SQuAD (link) 60k
German MLQA (Lewis et al., 2020)  5k
Hindi MLQA (Lewis et al., 2020)  5k
Vietnamese MLQA (Lewis et al., 2020)  5k
Preprocessing. To accommodate computational constraints, the analysis was limited to 2,000 question-answer (QA) pairs per language. A two-stage preprocessing pipeline was applied prior to sampling to ensure data quality and consistency: (1) duplicate removal, in which all redundant QA pairs were excluded; (2) answer validation, where pairs lacking valid responses, such as those with missing or ambiguous answers, were discarded.
4.2 Models
To investigate whether position bias in large language models (LLMs) arises from model-specific design and training choices or from language-specific characteristics, we evaluate five popular open-source multilingual models: Qwen2.5-7B-Instruct(Qwen et al., 2025), Llama3-8B-Instruct(Touvron et al., 2023), DeepSeek-7B-Chat(DeepSeek-AI, 2024), Gemma-7B-it(Team et al., 2024), and Mistral-7B-Instruct-v0.3(Jiang et al., 2023). All models support English, German, Russian, Vietnamese, and Hindi, yet differ in architecture and training paradigms. This selection enables us to disentangle model driven factors of position bias such as attention mechanisms and training data composition from language-related influences. It is important to note that DeepSeek-7B-Chat performed poorly on Hindi, therefore, we excluded its Hindi results from our analysis.
5 Evaluation
5.1 LLM as a Judge
Traditional statistical evaluation metrics of open-ended generations, such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and Meteor Universal (Denkowski & Lavie, 2014), are limited in their ability to differentiate between syntactically similar yet semantically distinct texts. Although semantic evaluation methods like BERTScore (Zhang et al., 2019) were developed to overcome these shortcomings, our experimental findings indicate that BERTScore exhibits insufficient discriminative power, frequently failing to capture subtle distinctions between correct and incorrect responses. Consequently, we employ the LLM as a judge framework (Zheng et al., 2023) and select Mistral Large [^2] https://mistral.ai/news/mistral-large-2407 as the evaluator for the following reasons: (1) prior research demonstrates its strong alignment with human judgments and generalizability across diverse tasks (Bavaresco et al., 2024; Kim et al., 2024); (2) it provides a freely accessible API for research, facilitating large-scale evaluation; (3) its architectural design differs from the majority of models used for response generation, thereby reducing potential bias toward self-generated outputs. The evaluator assesses question-answer pairs using a structured prompt that includes the question, ground truth, and model-generated answer. It assigns a label of $1$ for correct answers and 0 for incorrect ones; accuracy is adopted as the primary metric. Further details regarding the prompts and evaluation methodology are provided in Appendix F.
5.2 Human Evaluation
To validate the reliability of the large language model (LLM) as an evaluative judge, human annotation was conducted on a set of 150 questions in both English and Russian. The responses generated by the Llama model were annotated using the Overlap-3 metric, with domain experts adhering to the same evaluation criteria as the automated judge. Inter-annotator agreement was quantified using Krippendorff’s $\alpha$ (Krippendorff, 2011), yielding a mean $\alpha=0.755$, which indicates a high degree of assessment reliability. Further evaluation of alignment between the automated judge and human annotators is conducted by computing the Pearson correlation coefficient $r$ between the judge’s scores and the majority vote derived from human annotations. A strong mean correlation of $r=0.716$ was observed, indicating substantial agreement. Additional details regarding the annotation procedure are provided in Appendix F.
6 Experiments and Results
6.1 Sanity Check for evaluation procedure
Before proceeding with the remaining experiments, we first perform a simple sanity check to demonstrate that without any relevant context model performance drops significantly on two benchmarks: SQuAD2.0 (English) and MTS-SQuAD (Russian) with Llama3.1 8B model. In these experiment we wanted to answer a question if the model can generate accurate responses based solely on its internal knowledge. Additionally, we measure predictive entropy and show that it increases when no relevant context is provided. While these results are neither novel nor surprising, they serve to validate our overall evaluation procedure. The results are presented in Figure 2.
Figure 2: Results in the presence and absence of contextual information for Llama3.1 8B model on SQuAD2.0 (en) and MTS-SQuAD (ru). We use a single relevant context to evaluate whether the provision of such information enhances model performance or if the model can generate accurate responses based solely on its internal knowledge.
6.2 Overall Results Across Languages and Models
The second step of our evaluation assesses overall model performance and dataset difficulty. Tables 3 and 3 show performance averaged across all positional configurations and prompting strategies (full results are provided in Appendix H). Language performance is highest for English (SQuAD2.0: 0.896) and Russian (MT-SQuAD: 0.865), but drops for German (0.628) and especially Hindi and Vietnamese ($\approx$0.47), indicating greater difficulty in lower-resource or morphologically complex languages. For model performance, Qwen2.5-7B leads (0.754), followed by Mistral-7B (0.743), while DeepSeek-7B (0.711) and Llama3.1-8B (0.690) perform moderately. Gemma-7B lags substantially (0.486), suggesting weaker multilingual capabilities or architectural limitations.
Table 2: Performance on specific language averaged by models, positions and strategies
Dataset Accuracy
SQuAD2.0 (en) 0.896
MTS-SQuAD (ru) 0.865
MLQA (de) 0.628
MLQA (hi) 0.471
MLQA (vi) 0.475
Table 3: Performance of specific model averaged by datasets, positions, and strategies
Model Accuracy
DeepSeek 7B 0.711
Gemma 7B 0.486
Mistral 7B 0.743
Llama3.1 8B 0.690
Qwen2.5 7B 0.754
6.3 Position bias is mostly driven by models
While the “Lost in the Middle” phenomenon holds true overall, all models perform worst in the middle, as shown in Table 5, preferences for the beginning versus the end vary across datasets and models. We observe both model-specific bias (Table 6.3) and language-specific bias (Table 6.3). However, the aggregated results in Table 6 suggest that positional bias is primarily model-driven, as discrepancies across positions are more pronounced between models than across languages. Specifically, DeepSeek-7B, Mistral-7B, and Qwen2.5-7B exhibit a preference for later positions, whereas Llama3.1-8B and Gemma-7B favor earlier ones. The strongest positional effect is observed for DeepSeek-7B and Llama3.1-8B, with a performance difference of approximately 0.1 points between positions.
Table 4: Overall performance across strategies, the worst performance is highlighted in red.
Metric Aligned All Zero No Scores
Accuracy 0.695 0.610 0.721
Table 5: Overall performance across positions, the worst performance is highlighted in red.
Metric TOP MIDDLE BOTTOM
Accuracy 0.684 0.666 0.676
Table 6: Model-wise and Langue-wise position bias, highest accuracies reflect bias and are highlighted in bold-green. Results represent accuracy averaged across all strategies Aligned, All Zero and No Scores presented in Table 7 and Table 8. Position indicates correct context position among all contexts.
Position Dataset Mean accuracy Model Mean accuracy
TOP SQuAD2.0 (en) 0.904 DeepSeek 7B 0.683
MIDDLE 0.894 0.697
BOTTOM 0.891 0.753
TOP MTS-SQuAD (ru) 0.861 Gemma 7B 0.497
MIDDLE 0.861 0.492
BOTTOM 0.873 0.470
TOP MLQA (de) 0.646 Mistral 7B 0.748
MIDDLE 0.615 0.729
BOTTOM 0.624 0.751
TOP MLQA (hi) 0.493 Llama3.1 8B 0.743
MIDDLE 0.452 0.671
BOTTOM 0.469 0.655
TOP MLQA (vi) 0.480 Qwen2.5 7B 0.752
MIDDLE 0.463 0.744
BOTTOM 0.483 0.767
Table 7: Language specific bias with resulting accuracy averaged by models
Dataset Position Aligned All Zero No Scores
SQuAD2.0 (en) TOP 0.918 0.874 0.921
MIDDLE 0.919 0.844 0.920
BOTTOM 0.916 0.835 0.922
MTS-SQuAD (ru) TOP 0.877 0.823 0.883
MIDDLE 0.873 0.825 0.883
BOTTOM 0.889 0.841 0.888
MLQA (de) TOP 0.674 0.579 0.684
MIDDLE 0.658 0.525 0.661
BOTTOM 0.664 0.542 0.666
MLQA (hi) TOP 0.518 0.415 0.544
MIDDLE 0.483 0.353 0.520
BOTTOM 0.519 0.363 0.525
MLQA (vi) TOP 0.466 0.410 0.565
MIDDLE 0.452 0.379 0.558
BOTTOM 0.486 0.397 0.566
Table 8: Model specific bias with resulting accuracy averaged by datasets
Model Position Aligned All Zero No Scores
DeepSeek 7B TOP 0.683 0.669 0.696
MIDDLE 0.702 0.684 0.704
BOTTOM 0.761 0.743 0.756
Gemma 7B TOP 0.507 0.507 0.573
MIDDLE 0.515 0.515 0.571
BOTTOM 0.485 0.485 0.556
Mistral 7B TOP 0.755 0.709 0.779
MIDDLE 0.743 0.684 0.761
BOTTOM 0.762 0.716 0.776
Llama3.1 8B TOP 0.772 0.648 0.808
MIDDLE 0.720 0.517 0.777
BOTTOM 0.735 0.481 0.748
Qwen2.5 7B TOP 0.768 0.715 0.772
MIDDLE 0.750 0.715 0.767
BOTTOM 0.780 0.745 0.777
6.4 Positional guidance
Sensitivity to prompt guidance. For all considered languages incorrect relevance scoring leads to significant performance decrease. It varies from $4.3\%$ drop for English (0.918 Aligned vs 0.874 All Zero) to $15.6\%$ for Hindi (0.519 Aligned vs 0.363 All Zero). Similar patterns of positional guidance instability are observed for models. All models except Gemma exhibit pronounced sensitivity to positional cues when contextual scoring is perturbed. The introduction of misleading scores results in a marked decline in accuracy ranging from $1.4\%$ drop for DeepSeek performance (from $0.683$ Aligned to $0.669$ All-Zero), to $25.4\%$ drop for Llama performance (from $0.735$ Aligned to $0.481$ All-Zero). The absence of fluctuations in accuracy for Gemma is probably observed, because of poor quality of this model for all setups and general inability to handle positional guidance. Overall, the majority of setups are sensitive to positional guidance, so such techniques could be utilized to handle position bias.
Score Omission Enhances Robustness. Notably, the No Scores consistently outperforms other strategies. For languages there is a little improvement for Russian in BOTTOM position( $0.889$ Aligned vs $0.888$ No Scores). Similarly, only for Qwen2.5 7B( $0.780$ Aligned vs $0.777$ No Scores) and DeepSeek 7B( $0.761$ Aligned vs $0.756$ No Scores) in BOTTOM position the minor accuracy improvements are observed. For all other setups the performance is deteriorate, ranging from $0.1\%$ to $10.2\%$ accuracy drop for languages and from $0.2\%$ to $7.1\%$ for models (Aligned vs No Scores). This performance degradation is particularly pronounced in low-resource languages such as Vietnamese. These findings challenge previous works (Zhang et al., 2024a) and force to thorough validation of guiding strategies.
Language-Specific Sensitivity. High-resource languages, such as English and Russian, demonstrated minimal performance variation across scenarios ($\Delta<2.5\%$ in English), whereas languages on MLQA benchmark exhibited more pronounced differences. A phenomenon potentially attributable to orthographic or syntactic properties that may mediate position bias.
6.5 Highest Entropy is not associated with the lowest performance.
While we observe that models perform most poorly in the middle, the highest predictive entropy is not always associated with the middle position (see Table 9). We speculate that this phenomenon may arise due to token homogenization and provide a formal analysis in Appendix D.
Table 9: Predictive Entropy across positions, models and strategies. Highest mean values across strategies are highlighted in red.
Model Position Aligned All Zero No Scores Entropy
DeepSeek 7B TOP 0.257 0.258 0.233 0.250
MIDDLE 0.250 0.254 0.227 0.243
BOTTOM 0.233 0.238 0.213 0.228
Gemma 7B TOP 0.189 0.189 0.194 0.191
MIDDLE 0.187 0.187 0.196 0.190
BOTTOM 0.189 0.189 0.200 0.193
Mistral 7B TOP 0.194 0.194 0.217 0.202
MIDDLE 0.202 0.205 0.221 0.209
BOTTOM 0.199 0.201 0.215 0.205
Llama3.1 8B TOP 0.248 0.258 0.217 0.241
MIDDLE 0.251 0.241 0.232 0.241
BOTTOM 0.254 0.238 0.241 0.244
Qwen2.5 7B TOP 0.101 0.097 0.105 0.101
MIDDLE 0.106 0.104 0.112 0.107
BOTTOM 0.106 0.101 0.108 0.105
6.6 Word Order Analysis
We additionally perform investigation if there is the relationship between the position of relevant context, model behavior, and the dominant word order of a language. Since we find no evidence to suggest that position bias influences models to favor specific word orders we discuss our methodology and detailed results in Appendix E.
7 Conclusion
This study reveals that position bias in multilingual LLMs is primarily model-driven, contradicting the assumed universal early-token preference—with architectures like Qwen2.5-7B and DeepSeek-7B favoring late positions, while Llama-3.1-8B prefers early ones. Language-specific effects exist but are secondary. Surprisingly, explicitly signaling context relevance via prompt-based relevance scores consistently harms performance, especially in low-resource languages. Moreover, the worst accuracy (when relevant context is in the middle) does not correspond to higher output entropy, indicating models are confidently wrong under positional disadvantage. These findings challenge common RAG and prompting practices and underscore the need for model- and language-aware context handling.
8 Limitations
Since LLM-as-a-Judge was utilized to assess the correctness of open-ended question-answering task, our methodology depends on its performance critically.
Our evaluation used 2,000 question–answer pairs per language. Across nine experimental conditions and five models, this amounts to $9\times 5\times 2000=70000$ model evaluations per language, a computationally intensive effort. Given this scale, we took care to ensure statistical rigor. Specifically, we performed pairwise t-tests with Holm–Bonferroni correction for the three positional comparisons: (1) top vs. middle, (2) top vs. bottom, and (3) middle vs. bottom. For every dataset and model, at least one of these (two on average) comparisons yielded statistically significant differences in accuracy (p < 0.05). However, when aggregating results across models, we found no significant differences between languages. The same pattern held for predictive entropy values.
Acknowledgment on LLM assisted writing:
This paper used open access Qwen3-Max, in some parts of the paper, for proofreading and text rephrasing in accordance with formal style.
Appendix
Appendix A RAG systems with context reordering
REAR (Wang et al., 2024a) – integrates document relevance scores into LLMs via embeddings, guiding generation to use internal knowledge (low relevance) or external evidence (high relevance).
Long-Context LLMs Meet RAG (Jin et al., 2024) – addresses the "lost-in-the-middle" problem by reordering retrieved documents, placing highest-scoring ones at sequence boundaries to optimize LLM attention.
OP-RAG (Yu et al., 2024a) – order preserving RAG preserves original document order (vs. sorting chunks), demonstrating improved answer quality through position-aware context organization. However, authors do not mention multidocument scenario.
Clustering & Reordering RAG (Alessio et al., 2024) - cluster sentences by query similarity and sort clusters in descending similarity order for improved retrieval quality.
Appendix B Datasets
SQuAD2.0333https://huggingface.co/datasets/rajpurkar/squad_v2 (Rajpurkar et al., 2018) is an English reading-comprehension benchmark built on Wikipedia passages. It combines 100 000 span-answerable questions from SQuAD 1.1 (Rajpurkar et al., 2016) with more than 50 000 adversarial questions whose answers are deliberately absent.
MTS-SQuAD [^4] https://huggingface.co/datasets/MTS-AI-SearchSkill/MTSBerquad is an extension of the SberQuAD dataset (Efimov et al., 2020) which is a Russian counterpart of SQuAD 2.0. It includes more than 60 000 quesion-answer pairs with improved readability and consistency.
MLQA [^5] https://github.com/facebookresearch/MLQA (Lewis et al., 2020) is a multilingual benchmark built from aligned Wikipedia passages including 12 000 question-answer pairs in English and about 5 000 in each of the other six languages: Arabic, German, Spanish, Hindi, Vietnamese, and Simplified Chinese. Among these languages, we utilized German, Hindi and Vietnamese.
Appendix C Technical Details
Models inference. To achieve reproducibility of the obtained results, the LLM-inference was performed using a deterministic generation strategy. The following hyperparameter were used/set: "max_new_tokens" – 1024, "do_sample" – False, "num_beams" – 1.
Computational Resources. The experiments were run in a Docker container on a dedicated server with the following hardware: CPU: AMD Ryzen 9 7900X 12-Core Processor, GPU: NVIDIA GeForce RTX 3090 24GB, RAM: Kingston FURY Beast Black 32GB, SSD: M.2 NVMe Samsung 990 PRO 1T.
Required GPU-time for experiments. In total it is required approximately $50$ GPU-hours to reproduce the experiments.
Appendix D Effect of position bias on Entropy
This section formalizes the propagation of position bias toward initial tokens across layers in transformer-based large language models (LLMs) and examines its effect on attention entropy. Under the assumption of a standard multi-head self-attention architecture, we derive conditions under which such bias leads to homogenization of token representations, consequently increasing entropy in the final layer.
Notation. Let $X^{(0)}=[x_{1},x_{2},\dots,x_{n}]\in\mathbb{R}^{d\times n}$ denote the input token embeddings at layer 0, where $x_{1}$ is the first token and $d$ is the embedding dimension. At each layer $l\geq 1$, the self-attention operation computes:
$A^{(t)}=\text{softmax}\left(X^{(l)}W_{Q}^{(l)}(X^{(l)}W_{K}^{(t)})^{\top}\right)$ (2)
$X^{(t+1)}=W_{O}^{(l)}A^{(l)}X^{(l)}W_{V}^{(l)}$ (3)
where $W_{Q}^{(l)},W_{K}^{(l)},W_{V}^{(l)},W_{O}^{(l)}\in\mathbb{R}^{d\times d}$ are learnable projection matrices, $A^{(l)}\in\mathbb{R}^{n\times n}$ contains the attention weights and ${\sqrt{d_{QK}}}=1$ for simplicity.
Assumptions. To isolate the effect of position bias, we make simplifying assumptions: (1) dominant first token attention, (2) position bias does not change over layers and (3) attention $A$ can be represented as a linear combination of contextual attention $A^{\mathrm{cont}}$ and positional attention $A^{\mathrm{pos}}$. For all layers $l\geq 1$, the positional attention weights $A_{i,j}^{pos_{(l)}}$ are sharply concentrated on the first token:
$A_{i,j}^{pos_{(l)}}\approx\begin{cases}1&\text{if }j=1,\\ 0&\text{otherwise}.\end{cases}$
Or in vector form $A^{pos_{(l)}}\approx\mathbf{1}_{i}^{\top}\mathbf{e}_{1}$. Linear combination of attention: $A=\lambda_{1}\,{A}^{\text{cont}}+\lambda_{2}\,{A}^{\text{pos}},\qquad\lambda_{1}+\lambda_{2}=1;\lambda_{1},\lambda_{2}\in[0;1]$ where $\lambda$ are normalizing weights for each attention type.
Token Homogenization. Under these assumptions, the hidden state of token $i$ at layer $l$ becomes:
$x_{i}^{(l)}$ $\approx W_{O}^{(l)}\cdot\left[\sum_{j=1}^{n}\left(\lambda_{1}A_{i,j}^{\text{pos}(l)}+\lambda_{2}A_{i,j}^{\text{cont}(l)}\right)W_{V}^{(l)}x_{j}^{(l-1)}\right]$ (4)
$=\lambda_{2}W_{O}^{(l)}W_{V}^{(l)}\cdot\mathbf{1}_{i}^{\top}\mathbf{e}_{1}x_{i}^{(l-1)}+\lambda_{1}x_{i}^{\text{con}(l)}$
$=\lambda_{2}P^{(l)}x_{i}^{(l-1)}+\lambda_{1}x_{i}^{\text{con}(l)},\forall i.$
Where $P^{(l)}=W_{O}^{(l)}W_{V}^{(l)}$. If $\lambda_{2}>\lambda_{1}$ recursively applying this across $L$ layers yields:
$x_{i}^{(L)}\approx x_{1}^{(0)}\quad\forall i,$ (5)
implying all tokens collapse to a copy of the initial first token embedding $x_{1}^{(0)}$ (up to projection transformations). Although token collapse would not happen in a real scenario, for example, due to residual connections, tokens may become more similar to tokens under position bias.
Entropy Dynamics. Let $H_{A}^{(l)}$ denote the general attention entropy at layer $l$. As tokens homogenize ($x_{i}^{(l)}\approx x_{1}^{(0)}$), queries and keys become indistinguishable, causing contextual attention weights to approach uniformity: $A_{i,j}^{\text{con}_{(l)}}\approx\frac{1}{n}\quad\forall i,j$. This results in the maximization of entropy: $H_{A}^{(l)}\to\log n$.
Aligned positional and contextual attention. This section establishes a theoretical connection between entropy, attention mechanisms, and position bias under several modeling assumptions. We demonstrate that alignment between contextual attention $A^{\text{con}}$ and positional attention $A^{\text{pos}}$ increases the likelihood of homogenization. This leads to a counterintuitive outcome: when the relevant context coincides with the model’s inherent positional bias, the model allocates greater attention to the corresponding tokens, ultimately elevating entropy. Further empirical support for these findings is provided in Section D, where we observe that minimal entropy does not consistently coincide with alignment between contextual relevance and positional bias.
Predictive and Attention Entropy. Attention mechanisms are designed to prioritize relevant tokens within a sequence. When attention is uniformly distributed (indicating high entropy), the model is unable to effectively leverage contextual cues. This lack of focused attention results in diminished predictive signals, thereby increasing uncertainty in the model’s output predictions (manifested as higher predictive entropy). [^6] For example, in the sentence ”The capital of France is ***,” the token ”France” is critical for accurate prediction. If attention is uniform, the model may assign equal weight to less relevant tokens (e.g., ”The” or ”of”), failing to emphasize ”France.” This ambiguity elevates uncertainty in predicting the subsequent token, such as ”Paris.”
Multilingual Caveat. In multilingual large language models (LLMs), position bias may vary across different layers; for instance, early layers often prioritize English tokens, whereas later layers tend to align more closely with the language of the input prompt (Zhong et al., 2024; Schut et al., 2025). This shift may introduce language-specific positional preferences, thereby challenging the assumption of static position bias. Nevertheless, if homogenization occurs, the overall entropy dynamics remain consistent.
Appendix E Word Order Analysis
In this section, we perform additional investigation if there is the relationship between the position of relevant context, model behavior, and the dominant word order of a language. Specifically, we examine whether position bias amplifies or diminishes the influence of a model’s dominant language word order. Our analysis focuses on the No Scores configuration with five contexts, using Hindi and German as representative languages due to their non-SVO dominant word orders, as documented in The World Atlas of Language Structures. Sentences are parsed using Stanza (Qi et al., 2020), following a methodology similar to that of (Choi et al., 2021). For each verb, we identify its dependents; if a verb has both a subject (indicated by a "subj" relation, or such a relation for the nearest preceding verb connected via a "conj" dependency) and an object (indicated by an "obj" relation), we record the word order of these triplets using the abbreviations "S", "V", and "O". The distribution of "SVO" and "SOV" patterns is then analyzed relative to all extracted triplets.
We find no evidence to suggest that position bias influences models to favor specific word orders. For instance, in Hindi a predominantly subject–object–verb (SOV) language—one might expect subject–verb–object (SVO) rates to be lowest when relevant context appears last for Qwen and first for Llama, however, this pattern is not observed (Table 10). In German, where dominant word order varies by clause type (SVO in main clauses and SOV in subordinate clauses;  (Dryer, 2013)), we examine the prevalence of SVO (quantified as the SVO–SOV difference) in Table 11. Here, also, no association with position bias is detected, as the observed prevalence appears to arise naturally from the proportion of complex sentences.
Table 10: Percentage of word orders for Hindi, cells where context placements align with position bias are highlighted in green. Bolded results indicate stronger alignment with expected word order.
Model Word order TOP MIDDLE BOTTOM
Llama3.1 8B SVO 5.39 5.12 4.75
SOV 88.52 88.49 88.43
Qwen2.5 7B SVO 2.34 4.15 2.96
SOV 93.98 91 91.78
Table 11: Prevalence of SVO over SOV word order and the proportion of complex sentences in German
Model Position Difference SVO - SOV % of Complex Sentences
Llama3.1 8B TOP 26.88 13.56
MIDDLE 32.13 12.13
BOTTOM 40.04 10.88
Qwen2.5 7B TOP 34.17 8.59
MIDDLE 29.56 9.54
BOTTOM 29.78 8.96
For model-level analysis, the following distinctions are observed: (1) Llama generates a higher proportion of subject–verb–object (SVO) sentences than Qwen across Hindi, German, and Russian (Table 12)—note that Russian is included despite its typological preference for SVO order due to its rich morphological system, which permits considerable word-order flexibility (Ghosh et al., 2024a); (2) Llama produces syntactically more complex sentences than Qwen across all three contextual positions (Table 13); (3) both models exhibit a tendency to generate more complex sentences when relevant context is provided in the initial position (Table 14).
Table 12: Percentage of SVO structures in generated responses across all three positions of relevant context
Model Hindi German Russian
Llama3.1 8B 4.8 - 5.4% 60.3 - 67.5% 95.9 - 96.2%
Qwen2.5 7B 2.3 - 4.2% 57.9 - 61.4% 93.6 - 94.5%
Table 13: Mean percentage of complex sentences among all sentences containing at least one triplet of a subject, verb, and object
Model Hindi German Russian
Llama3.1 8B 9.94 % 12.19 % 7.19%
Qwen2.5 7B 6.88% 9.03 % 7%
Table 14: The context position in which the proportion of complex sentences is highest, given a specific model and language
Language Llama3.1 8B Qwen2.5 7B
TOP MIDDLE BOTTOM TOP MIDDLE BOTTOM
Hindi $\checkmark$ $\checkmark$
German $\checkmark$ $\checkmark$
Russian $\checkmark$ $\checkmark$
Appendix F Evaluation Details
F.1 LLM-as-a-Judge Verification
Krippendorff’s alpha and Pearson correlation coefficients, calculated for each experimental setup can be seen in Tables 15 and 16 correspondingly. Comparison of human and Llama evaluation can be seen in Figure 3.
Table 15: Krippendorff’s alpha coefficient, calculated for each experimental setup
Language Position Aligned All Zero No Scores Mean
English TOP 0.783 0.663 0.595 0.726
MIDDLE 0.611 0.861 0.685
BOTTOM 0.704 0.916 0.718
Russian TOP 0.814 0.825 0.674 0.783
MIDDLE 0.742 0.865 0.695
BOTTOM 0.801 0.855 0.776
Table 16: Pearson correlation coefficient, calculated for each experimental setup
Language Position Aligned All Zero No Scores Mean
English TOP 0.612 0.632 0.604 0.727
MIDDLE 0.739 0.83 0.791
BOTTOM 0.738 0.908 0.685
Russian TOP 0.488 0.704 0.518 0.705
MIDDLE 0.709 0.87 0.769
BOTTOM 0.669 0.888 0.732
F.2 Human and LLM instructions
We prompt Mistral-Large to judge whether the LLM responses are correctly answering questions. For each dataset we create an evaluation prompt on the language of this dataset and add 4 shots as examples of judgments. The resulting prompt consists of system prompt "You are an AI assistant who speaks English.", which we translate to other languages and user prompt.
For human annotation we consider only English and Russian languages, since our annotators speaks these languages. We use the same instructions as for LLM-as-a-Judge settings, omitting shots.
F.3 Human Annotators Information
Annotation was conducted by the authors of the work, so no additional recruitment or payment are required on this stage. All assessors held bachelor’s degree and had prior experience in the evaluation of LLM responses.
Figure 3: Human evaluation and LLM as a Judge for Correct Scores(Aligned), All Zero, and No Scores strategies at three postions (TOP, MIDDLE, BOTTOM) for two languages: (a) English and (b) Russian. Bars in green represent human evaluations, while the blue bars represent the Llama model.
Appendix G Context Volume Analysis
From Figure 4 we can observe that our Aligned strategy does not have an effect on position bias with increasing of information load. With context quantity $N=15$ for Llama3.1-8B we can see significant accuracy decrease, compared to other quantities. From the other hand, for Qwen2.5-7B position bias does not correlate with passed number of contexts. This result can be explained by the fact that for Qwen2.5-7B training larger dataset with long contexts was used, compared to Llama3.1-8B. This feature increase for Qwen2.5-7B the size of its attention window and allowed to conditioning on a larger amount of input knowledge during response generation.
Figure 4: Accuracy dependence on the number of contexts, added to the user-prompt, and position of the relevant context in a list with Aligned placement strategy: (a) Llama3.1-8B; (b) Qwen2.5-7B
Appendix H Non Aggregated results for main experiments
Our non aggregated results for LLM as an accuracy and entropy are presented in Tables 17,18 and 19, 20 correspondingly.
Table 17: Accuracy for three models: DeepSeek 7B Chat, Gemma 7B and Mistral 7B
DeepSeek 7B Chat Gemma 7B Mistral 7B
Dataset Position Aligned All Zero No Scores Aligned All Zero No Scores Aligned All Zero No Scores
TOP 0.907 0.877 0.913 0.815 0.683 0.831 0.970 0.956 0.961
MIDDLE 0.922 0.887 0.921 0.849 0.614 0.832 0.958 0.929 0.957
Squadv2 (en) BOTTOM 0.953 0.929 0.950 0.782 0.534 0.815 0.963 0.939 0.958
TOP 0.804 0.791 0.841 0.775 0.655 0.797 0.939 0.926 0.936
MIDDLE 0.834 0.829 0.843 0.777 0.701 0.819 0.920 0.911 0.924
MTS-SQuAD (ru) BOTTOM 0.884 0.875 0.881 0.782 0.707 0.824 0.932 0.928 0.930
TOP 0.737 0.732 0.720 0.477 0.324 0.524 0.829 0.799 0.813
MIDDLE 0.735 0.708 0.724 0.508 0.289 0.486 0.815 0.750 0.801
MLQA (de) BOTTOM 0.799 0.772 0.777 0.432 0.244 0.441 0.823 0.787 0.819
TOP 0.431 0.372 0.229 0.520 0.439 0.601
MIDDLE 0.402 0.334 0.237 0.505 0.425 0.558
MLQA (hi) BOTTOM - 0.393 0.349 0.228 0.546 0.471 0.585
TOP 0.284 0.277 0.309 0.039 0.021 0.484 0.520 0.428 0.584
MIDDLE 0.317 0.314 0.330 0.039 0.020 0.483 0.517 0.406 0.566
MLQA (vi) BOTTOM 0.407 0.397 0.415 0.035 0.014 0.471 0.545 0.457 0.588
Table 18: Accuracy for two models: Llama3.1 8B and Qwen2.5 7B
Dataset Position Llama3.1 8B Qwen2.5 7B
Aligned All Zero No Scores Aligned All Zero No Scores
SQuAD2.0 (en) TOP 0.955 0.929 0.951 0.943 0.927 0.952
MIDDLE 0.926 0.858 0.946 0.941 0.931 0.945
BOTTOM 0.932 0.826 0.930 0.949 0.948 0.956
MTS-SQuAD (ru) TOP 0.928 0.837 0.906 0.938 0.909 0.934
MIDDLE 0.908 0.768 0.888 0.927 0.917 0.944
BOTTOM 0.911 0.761 0.866 0.937 0.935 0.939
MLQA (de) TOP 0.680 0.485 0.719 0.648 0.553 0.644
MIDDLE 0.602 0.312 0.668 0.629 0.566 0.628
BOTTOM 0.613 0.307 0.646 0.655 0.602 0.649
MLQA (hi) TOP 0.532 0.309 0.729 0.591 0.540 0.619
MIDDLE 0.448 0.112 0.676 0.579 0.543 0.609
BOTTOM 0.494 0.038 0.644 0.642 0.592 0.643
MLQA (vi) TOP 0.764 0.679 0.737 0.722 0.644 0.713
MIDDLE 0.714 0.536 0.707 0.676 0.619 0.707
BOTTOM 0.726 0.473 0.656 0.718 0.647 0.700
Table 19: Predictive entropy values for three models: DeepSeek 7B Chat, Gemma 7B and Mistral 7B
DeepSeek 7B Chat Gemma 7B Mistral 7B
Dataset Position Aligned All Zero No Scores Aligned All Zero No Scores Aligned All Zero No Scores
TOP 0.140 0.148 0.121 0.131 0.160 0.116 0.122 0.124 0.126
MIDDLE 0.132 0.137 0.116 0.141 0.166 0.110 0.125 0.130 0.131
SQuAD2.0 (en) BOTTOM 0.128 0.129 0.114 0.146 0.175 0.113 0.117 0.122 0.125
TOP 0.225 0.222 0.240 0.183 0.229 0.155 0.185 0.191 0.204
MIDDLE 0.216 0.222 0.232 0.164 0.204 0.149 0.193 0.200 0.210
MTS-SQuAD (ru) BOTTOM 0.208 0.212 0.231 0.167 0.199 0.145 0.190 0.197 0.207
TOP 0.237 0.241 0.240 0.234 0.228 0.219 0.276 0.288 0.292
MIDDLE 0.238 0.249 0.238 0.218 0.209 0.225 0.286 0.321 0.295
MLQA (de) BOTTOM 0.222 0.233 0.228 0.215 0.197 0.235 0.274 0.300 0.287
TOP 0.279 0.325 0.366 0.213 0.202 0.238
MIDDLE 0.302 0.349 0.382 0.221 0.204 0.236
MLQA (hi) BOTTOM - 0.312 0.343 0.391 0.227 0.214 0.237
TOP 0.427 0.423 0.332 0.120 0.112 0.115 0.175 0.166 0.225
MIDDLE 0.412 0.409 0.322 0.110 0.106 0.113 0.185 0.171 0.232
MLQA (vi) BOTTOM 0.374 0.379 0.277 0.104 0.104 0.114 0.185 0.171 0.218
Table 20: Predictive entropy values for two models Llama3.1 8B and Qwen2.5 7B
Dataset Position Llama3.1 8B Qwen2.5 7B
Aligned All Zero No Scores Aligned All Zero No Scores
SQuAD2.0 (en) TOP 0.232 0.248 0.194 0.092 0.093 0.100
MIDDLE 0.237 0.239 0.201 0.094 0.102 0.100
BOTTOM 0.241 0.240 0.206 0.093 0.099 0.097
MTS-SQuAD (ru) TOP 0.203 0.224 0.169 0.105 0.111 0.129
MIDDLE 0.217 0.219 0.182 0.112 0.118 0.126
BOTTOM 0.214 0.222 0.191 0.112 0.109 0.126
MLQA (de) TOP 0.231 0.204 0.220 0.117 0.106 0.115
MIDDLE 0.220 0.158 0.225 0.126 0.118 0.130
BOTTOM 0.225 0.152 0.220 0.128 0.118 0.125
MLQA (hi) TOP 0.303 0.320 0.237 0.078 0.075 0.076
MIDDLE 0.294 0.290 0.257 0.082 0.077 0.082
BOTTOM 0.289 0.277 0.275 0.083 0.082 0.080
MLQA (vi) TOP 0.272 0.296 0.267 0.114 0.101 0.107
MIDDLE 0.288 0.299 0.295 0.117 0.103 0.120
BOTTOM 0.298 0.298 0.313 0.111 0.100 0.112
Appendix I Prompts
Our user–prompts for LLM-inference in terms of context placement strategies can be seen in Tables 21, 22. For Aligned and All Zero strategies items in a contexts–list has a following format: "- [{score}] {document}". For No Scores strategy items has the following format: "- {document}". As a system–prompt the same instruction was used for all languages (translated correspondingly): "You are an AI assistant who helps solve user issues.".
Table 21: User–prompts in five languages for LLM–inference in the No Scores context placement strategy
Language User Prompt
English [Uncaptioned image]
Russian [Uncaptioned image]
German [Uncaptioned image]
Hindi [Uncaptioned image]
Vietnamese [Uncaptioned image]
Table 22: User–prompts in five languages for LLM–inference in the Aligned and All Zero context placement strategies
Language User Prompt
English [Uncaptioned image]
Russian [Uncaptioned image]
German [Uncaptioned image]
Hindi [Uncaptioned image]
Vietnamese [Uncaptioned image]